Max is 24 years older than his son, Liam. In two years, Liam will be half as old as Max. How old is Max now? _____yeadr old
Answer by jorel555(1290)

Let L be Liam. Then Max is L + 24. So:
L+2=1/2(L+24+2)
2L+4=L+26
L=22
Liam is 22, Max is 46. ☺☺☺☺